# A Deep Automotive Radar Detector using the RaDelft Dataset

### v1

@@ Abstract
The detection of multiple extended targets in complex environments using high-resolution automotive radar is considered. A data-driven approach is proposed where unlabeled synchronized lidar data is used as ground truth to train a neural network with only radar data as input. To this end, the novel, large-scale, real-life, and multi-sensor RaDelft dataset has been recorded using a demonstrator vehicle in different locations in the city of Delft. The dataset, as well as the documentation and example code, is publicly available for those researchers in the field of automotive radar or machine perception.
The proposed data-driven detector is able to generate lidar-like point clouds using only radar data from a high-resolution system, which preserves the shape and size of extended targets. The results are compared against conventional CFAR detectors as well as variations of the method to emulate the available approaches in the literature, using the probability of detection, the probability of false alarm, and the Chamfer distance as performance metrics. Moreover, an ablation study was carried out to assess the impact of Doppler and temporal information on detection performance. The proposed method outperforms the different baselines in terms of Chamfer distance, achieving a reduction of 75% against conventional CFAR detectors and 10% against the modified state-of-the-art deep learning based approaches.
@@ II-A Terminology
  • Point cloud refers to a set of $\textrm{N}_{p}$ points, each containing L features that result from selecting only those cells containing ones in a 3D occupancy grid and converting them to Cartesian coordinates. For radar point clouds it is typically assumed that $\textrm{L}=5$, adding Doppler and power information to the three spatial dimensions, while for lidar point clouds $\textrm{L}=4$ since Doppler is not provided.
@@ II-B Radar Processing Pipeline Review
    The detection stage then identifies the cells that contain the targets. Usually, a combination of a CFAR detector in some dimensions and peak finding in the rest is used, though some works have also explored using machine learning algorithms [5, 6, 7, 8, 9, 10, 11]. In this stage, the data is often sparse since most of the space in the field of view does not reflect sufficient power or is simply empty. The detector outputs a 3D occupancy grid, but a conversion to point cloud is usually performed since it is a convenient format for visualizations or for dataset storage.
@@ III Related Work
III-A Radar Datasets Related Work
@@
III-B Radar Detectors Related Work
@@
Also, DL techniques have been applied to the radar detection problem [5, 6, 7, 8, 9, 10]. In [5], a DL detector is proposed, outperforming several 2D CA-CFAR detectors, but only tested in simulated data. In [41] and [8], the authors propose a similar network structure using 3 autoencoders in three 2D projections (range-angle, range-Doppler, and angle-Doppler) using the annotated dataset in [14] by a camera, avoiding full 3D detection. However, using camera detections as ground truth may be limited due to the 2D nature of camera images. On the other hand, the works [10] and [7] propose two different NNs, but both use the lidar point cloud as ground truth. Since lidar provides high-resolution 3D point clouds, it seems a more reasonable choice to serve as ground truth. The proposed method in [7] uses a neural network to detect targets only in the range-Doppler dimensions, followed by the angle estimation and a spatio-temporal filter to enhance the resulting point cloud. On the other hand, in [10], a novel sparse approach to use an NN to detect in the range-azimuth-elevation space is presented. However, the Doppler information is collapsed into a single value, preventing the network from learning the possible angular estimation enhancement due to its relationship with Doppler [24, 25]. Moreover, only the top 10% power cells are used as input to the network, and therefore, a pre-detection step is used, which can potentially remove target cells. This may be critical in automotive scenarios, where the angular sidelobes of close-range targets may be even 20dB higher than weakly-reflecting distant targets such as pedestrians.
@@ IV RaDelft Dataset
The dataset was recorded with the demonstrator vehicle presented in [42] with an additional Texas Instrument MMWCAS-RF-EVM [43] imaging radar mounted on the roof at 1.5 meters from the ground. The details of the radar and the waveform used are provided in Section IV-A. The collection was performed driving in multiple real-life scenarios in the city of Delft with different scene characteristics, such as suburban, university campus, and Delft old-town locations. Four different camera frames are shown in Fig. 2 to illustrate the differences in the environments. The output of the following sensors was recorded: a RoboSense Ruby Plus Lidar (128 layers rotating lidar) and the imaging radar board installed on the roof, a video camera (1936 × 1216 px, $\sim$30 Hz) mounted behind the windshield, and the ego vehicle’s odometry (filtered combination of RTK GPS, IMU, and wheel odometry, $\sim$100 Hz). The sensor setup can be seen in Fig. 3. All sensors were jointly calibrated following [44] and time synchronized. With a 10 Hz frame rate, each scene contains around 2500 radar frames, adding to a total of 16975 frames.
@@ IV-A Radar Configuration and Processing
Maximum Unambiguous Velocity (without extension) (m/s) 2.93
Maximum Unambiguous Velocity (with extension) (m/s) 20.56
@@
In this work, both undesirable effects of TDMA are solved using the overlapped virtual antennas present in the radar system with the algorithms provided in [46]. However, it is important to take into account that the maximum unambiguous velocity extension only works when a single target is present in a range-Doppler cell. Therefore, if multiple targets are folded into the same Doppler bin, or there are targets in different angles at the same range-Doppler bin, the algorithm will not be able address the problem. Since this work does not aim to solve the Doppler ambiguity problem in TDMA, the aforementioned constraint is accepted as a limitation of the current commercial radar system. Nevertheless, it is assumed that making the ADC samples directly available in our dataset can be valuable for the research community, for example to apply more advanced approaches for Doppler/velocity ambiguity in TDMA in the future.
The angle estimation can be performed once the TDMA effects have been compensated. It is important to remember that the resulting virtual array is a very sparse URA with some structures. While other research works deal with this type of array, for instance by trying to fill/interpolate the missing elements or applying compressive sensing techniques [47, 22], the core of this work is not to improve the angular estimation with sparse arrays. Therefore, a very simple approach of zero-filling and FFT processing has been applied. However, due to the sparseness of the radar antenna array in the Z-direction (shown in Fig. 3), grating lobes and high-side lobes appear in elevation. To mitigate this problem, the field of view in elevation has been restricted to ±$\pm 20^{\circ}$ degrees, and the elevation value with the highest power has been selected and saved, discarding the rest. Also, the azimuth estimation has been restricted to ±$\pm 70^{\circ}$ for two reasons. First, the angular estimation performance outside this region is rather poor, as:
@@ V Proposed Data Driven Detector
The first step of the method is to adapt the lidar point cloud to serve as the ground truth. Since the lidar system used in this work is mechanically rotating, it provides $360^{\circ}$ coverage. Therefore, the first step is to crop this as to the same FoV of the radar, i.e., ±$\pm 20^{\circ}$ in elevation & ±$\pm 70^{\circ}$ in azimuth, and a maximum range of 50m. To illustrate this difference in the FoV, Fig. 5b shows the cropped lidar point cloud compared to the original point cloud in Fig. 5a. Moreover, removing all the lidar points from the road surface is essential as the road surface is hardly visible to the radars and could lead to noisy ground truth for the training process. The Patchwork++ algorithm is used to this end [48]. After removing the road surface points, the resulting lidar point cloud can be seen in Fig. 5c.
@@
Figure 4: Overview of the proposed data-driven detector. The steps to generate the 3D lidar occupancy grid are shown on the top row, which will be then used as ground truth for training the neural network. The radar signal processing pipeline is shown at the bottom of the figure, and is needed to generate the input data for the network. RDC stands for range-Doppler-channel (no angle estimation), and RAED stands for range-azimuth-elevation-Doppler [18].
@@
These two parts of the proposed network are applied to each of the three considered frames independently, as shown in Fig. 6, but the weights of the layers are shared, and the output is concatenated into a $\mathrm{T}\times\mathrm{R}\times\mathrm{A}\times\mathrm{E}\ (3\times 500\times 240\times 44)$ tensor. Finally, to take into account the temporal relationship between the three frames, a third module composed of six 3D convolutional layers is included (referred to as TemporalCoherenceNetwork in Fig. 6).
One of the key characteristics of the radar data is the scene sparsity. Of all the voxels in the generated 3D occupancy grid, only around 1% contain targets. Therefore, this must be considered when selecting the loss function for training the neural network. In this work, the Focal loss [51] is used for this purpose, which handles class imbalances in a similar way to the weighted cross-entropy loss, and adds an extra modulating factor to focus on the hard cases. This is especially interesting in radar data, since high RCS targets can be easily detected, but low RCS or far targets are more challenging. In terms of training-testing split, 90% of the data from five of the seven recorded scenarios have been used to train the network using Adam optimizer, leaving 10% for validation. The network was trained using the DelftBlue Supercomputer [52] from TU Delft. The remaining two recorded scenarios are used as a test set, i.e., with data completely new, unseen for the network.
@@ VI Results
Two main performance metrics are used to evaluate the results of the proposed neural network: the usual probability of detection and probability of false alarm metrics, and the Chamfer distance (CD). In both cases, the lidar data is used as a reference, either in the occupancy grid format for the Pd and Pfa computation, or in the point cloud format for the Chamfer distance. While different definitions are given for the Chamfer distance in the literature, in this work the following is used:
@@
  • No Doppler & Quantile: an approach similar to the one presented in [10], where only those power cells with values higher than the 0.9 quantile are kept and the rest is set to zero. Furthermore, the Doppler information is collapsed by taking the mean over the Doppler dimension. This is used to ’sparsify’ the data and speed up processing, with the risk of cutting out weakly reflecting targets. Since there is no Doppler data anymore, the DopplerEncoder subnetwork is removed from the general architecture of the proposed data-driven detector.
  • No Doppler: the proposed method without any predefined threshold on the data, but with the Doppler information collapsed by taking the mean of the data and thus without the DopplerEncoder subnetwork in the architecture. This is done to perform an ablation study on the importance of the Doppler domain features and analyze the effect of each block independently.
@@
  • OSCFAR: a 2D OS CFAR in range-angle, followed by a 1D OS CFAR in Doppler. While multiple different CFAR alternatives have been tested (i.e., different combinations of CA and OS CFAR detectors), only the best implementation is reported here for conciseness. An analysis with different variations has been presented in [18] for completeness. Following [53], in this paper the rank has been set to 0.75 times the number of training cells, and no guard cells have been used.
As it can be seen in Table III, the highest Pd is achieved by the proposed method at the cost of a higher Pfa. On the other hand, applying the quantile cut and removing the Doppler information similarly to [10] drastically reduces the Pd from 63.5% to 39.3%, and worsens the Chamfer distance from 1.644m to 1.826m, but the Pfa is also reduced. Looking at the results for the other versions, it can be seen that this drop in performance is mostly due to the removal of the Doppler information. Using only the quantile-based threshold may be a good trade-off since the Chamfer distance is reduced from 1.644m to 1.430m at the cost of reducing the Pd. However, this may be a suitable alternative, especially due to the computational cost reduction. Looking at the version without the Temporal Coherence subnetwork, which is trained on single frames, it can be seen how all the metrics are worse than in the baseline. Thus, including temporal information in the network is a good strategy to boost performance, with the only downside of increasing slightly the training time due to the extra layers. Finally, it can be seen that the conventional OS CFAR is the method that performs the worst, with a much higher Chamfer distance of 6.73m.
In order to have a fairer comparison against the conventional CFAR detector, a 2D version of the proposed method has been also evaluated by disregarding the elevation information, as this can only be estimated rather poorly due to the unfavourable design of the radar array. To this end, the proposed NN has been trained without elevation information, discarding the virtual channels in the Z-direction and, thus, treating it as a ULA in the azimuth direction. For completeness, the implementation with quantile based threshold has also been assessed in this new analysis. The results are shown in Table III under the ’No Elevation cases’ rows. For these tests, the Pd of the OS CFAR approach is increased to 11.5%, but the Pfa is also raised. This is mainly due to detections triggered in the adjacent angle bins of a target generating ”ring like” patterns due to side lobes, a phenomenon also mentioned in [10]. Both the proposed method and the proposed method with the quantile based threshold are shown to outperform the conventional OSCFAR in the three metrics.
@@
Proposed Method 63.5 2.86 1.644
No Doppler & Quantile 39.3 0.87 1.823
No Doppler 39.4 0.85 1.826
Quantile 45.5 0.98 1.430
No Time (single frame) 61.9 3.15 1.943

### v3

@@ Abstract
The detection of multiple extended targets in complex environments using high-resolution automotive radar is considered. A data-driven approach is proposed where unlabeled synchronized lidar data is used as ground truth to train a neural network with only radar data as input. To this end, the novel, large-scale, real-life, and multi-sensor RaDelft dataset has been recorded using a demonstrator vehicle in different locations in the city of Delft. The dataset, as well as the documentation and example code, is publicly available for those researchers in the field of automotive radar or machine perception. The proposed data-driven detector is able to generate lidar-like point clouds using only radar data from a high-resolution system, which preserves the shape and size of extended targets. The results are compared against conventional CFAR detectors as well as variations of the method to emulate the available approaches in the literature, using the probability of detection, the probability of false alarm, and the Chamfer distance as performance metrics. Moreover, an ablation study was carried out to assess the impact of Doppler and temporal information on detection performance. The proposed method outperforms the different baselines in terms of Chamfer distance, achieving a reduction of 77% against conventional CFAR detectors and 28% against the modified state-of-the-art deep learning based approaches.
@@ II-A Terminology
  • Point cloud refers to a set of $\textrm{N}_{p}$ points, each containing L features that result from selecting only those cells containing ones in a 3D occupancy grid and converting them to Cartesian coordinates. For radar point clouds is typically assumed that $\textrm{L}=5$, adding Doppler and power information to the three spatial dimensions, while for lidar point clouds $\textrm{L}=4$ since Doppler is not provided.
@@ II-B Radar Processing Pipeline Review
    The detection stage then identifies the cells that contain the targets. Usually, a combination of a CFAR detector in some dimensions and peak finding in the rest is used, though some works have also explored using machine learning algorithms [5, 7, 8, 9, 6, 10]. In this stage, the data is often sparse since most of the space in the field of view does not reflect sufficient power or is simply empty. The detector outputs a 3D occupancy grid, but a conversion to point cloud is usually performed since it is a convenient format for visualizations or for dataset storage.
@@ III Related Work
III-A Radar Datasets
@@
III-B Radar Detectors
@@
Also, DL techniques have been applied to the radar detection problem [5, 7, 8, 9, 6, 11]. In [5], a DL detector is proposed, outperforming several 2D CA-CFAR detectors, but only tested in simulated data. In [41] and [8], the authors propose a similar network structure using 3 autoencoders in three 2D projections (range-angle, range-Doppler, and angle-Doppler) using the annotated dataset in [14] by a camera, avoiding full 3D detection. However, using camera detections as ground truth may be limited due to the 2D nature of camera images. Also, in [11], the authors propose a DL-based detector using bird-eye view radar data, but focusing only on vehicle detection. On the other hand, the works [6] and [7] propose two different NNs, but both use the lidar point cloud as ground truth. Since lidar provides high-resolution 3D point clouds, it seems a more reasonable choice to serve as ground truth. The proposed method in [7] uses a neural network to detect targets only in the range-Doppler dimensions, followed by the angle estimation and a spatio-temporal filter to enhance the resulting point cloud. On the other hand, in [6], a novel sparse approach to use an NN to detect in the range-azimuth-elevation space is presented. However, the Doppler information is collapsed into a single value, preventing the network from learning the possible angular estimation enhancement due to its relationship with Doppler [24, 25]. Moreover, only the top 10% power cells are used as input to the network, and therefore, a pre-detection step is used, which can potentially remove target cells. This may be critical in automotive scenarios, where the angular sidelobes of close-range targets may be even 20dB higher than weakly-reflecting distant targets such as pedestrians.
@@ IV RaDelft Dataset
The dataset was recorded with the demonstrator vehicle presented in [42] with an additional Texas Instrument MMWCAS-RF-EVM [43] imaging radar mounted on the roof at 1.5 meters from the ground. The details of the radar and the waveform used are provided in Section IV-A. The collection was performed driving in multiple real-life scenarios in the city of Delft with different scene characteristics, such as suburban, university campus, and Delft old-town locations. Four different camera frames are shown in Fig. 2 to illustrate the differences in the environments. The output of the following sensors was recorded: a RoboSense Ruby Plus Lidar (128 layers rotating lidar, 10Hz) and the imaging radar board installed on the roof, a video camera (1936 × 1216 px, $\sim$30 Hz) mounted behind the windshield, and the ego vehicle’s odometry (filtered combination of RTK GPS, IMU, and wheel odometry, $\sim$100 Hz). The sensor setup can be seen in Fig. 3. All sensors were jointly calibrated following [44] and time synchronized. With a 10 Hz frame rate, each scene contains around 2500 radar frames, adding to a total of 16975 frames.
@@ IV-A Radar Configuration and Processing
Maximum Unambiguous Velocity (without extension) (m/s) 2.48
Maximum Unambiguous Velocity (with extension) (m/s) 17.36
@@
In this work, both undesirable effects of TDMA are solved by using the overlapped virtual antennas present in the radar system with the algorithms provided in [46]. However, it is important to take into account that the maximum unambiguous velocity extension only works when a single target is present in a range-Doppler cell. Therefore, if multiple targets are folded into the same Doppler bin, or there are targets in different angles at the same range-Doppler bin, the algorithm will not be able address the problem. Since this work does not aim to solve the Doppler ambiguity problem in TDMA, the aforementioned constraint is accepted as a limitation of the current commercial radar system. Nevertheless, it is assumed that making the ADC samples directly available in our dataset can be valuable for the research community, for example to apply more advanced approaches for Doppler/velocity ambiguity in TDMA in the future.
The angle estimation can be performed once the TDMA effects have been compensated. It is important to remember that the resulting virtual array is a very sparse URA with some structures. While other research works deal with this type of array, for instance by trying to fill/interpolate the missing elements or applying compressive sensing techniques [47, 22], the core of this work is not to improve the angular estimation with sparse arrays. Therefore, a very simple approach of zero-filling and FFT processing has been applied. However, due to the sparseness of the radar antenna array in the Z-direction (shown in Fig. 3), grating lobes and high-side lobes appear in elevation. To mitigate this problem, the field of view in elevation has been restricted to ±$\pm 15^{\circ}$ degrees, and the elevation value with the highest power has been selected and saved, discarding the rest. Also, the azimuth estimation has been restricted to ±$\pm 70^{\circ}$ for two reasons. First, the angular estimation performance outside this region is rather poor, as:
@@ V Proposed Data Driven Detector
The first step of the method is to adapt the lidar point cloud to serve as the ground truth. For each radar cube, the closest lidar point cloud in time is selected based on the timestamps for both radar and lidar data, assuming that a small error due to different start times may be present. Since the lidar system used in this work is mechanically rotating, it provides $360^{\circ}$ coverage. Therefore, the first step is to crop this as to the same FoV of the radar, i.e., ±$\pm 20^{\circ}$ in elevation & ±$\pm 70^{\circ}$ in azimuth, and a maximum range of 50m. To illustrate this difference in the FoV, Fig. 5b shows the cropped lidar point cloud compared to the original point cloud in Fig. 5a. Moreover, removing all the lidar points from the road surface is essential as the road surface is hardly visible to the radars and could lead to noisy ground truth for the training process. The Patchwork++ algorithm is used to this end [48]. After removing the road surface points, the resulting lidar point cloud can be seen in Fig. 5c.
@@
Figure 4: Overview of the proposed data-driven detector. The steps to generate the 3D lidar occupancy grid are shown on the top row, which will be then used as ground truth for training the neural network. The radar signal processing pipeline is shown at the bottom of the figure, and is needed to generate the input data for the network. RDC stands for range-Doppler-channel (no angle estimation), and RAED stands for range-azimuth-elevation-Doppler [18]. RoI stands for region of interest.
@@
These two parts of the proposed network are applied to each of the three considered frames independently, as shown in Fig. 6, but the weights of the layers are shared, and the output is concatenated into a $\mathrm{T}\times\mathrm{R}\times\mathrm{A}\times\mathrm{E}\ (3\times 500\times 240\times 44)$ tensor. Finally, to take into account the temporal relationship between the three frames, a third module composed of six 3D convolutional layers is included (referred to as TemporalCoherenceNetwork in Fig. 6). It is important to notice that even if the output is a 3D occupancy grid for each frame, the power information on each cell is not lost, since the indices of the detected cells from such grid can be used to retrieve the corresponding intensity information from the original RAED cube.
One of the key characteristics of the radar data is the scene sparsity. Of all the voxels in the generated 3D occupancy grid, only around 1% contain targets. Therefore, this must be considered when selecting the loss function for training the neural network. In this work, the Focal loss [51] is used for this purpose, which handles class imbalances in a similar way to the weighted cross-entropy loss, and adds an extra modulating factor to focus on the hard cases. The Focal loss [51] is defined as:
$FL(p_{t})=-\alpha_{t}(1-p_{t})^{\gamma}\log(p_{t}),$ (4)
with
$p_{t}=\begin{cases}p&\textrm{if}\quad y=1\\ 1-p&\textrm{otherwise}\\ \end{cases}$ (5)
where $y\in\{\pm 1\}$ is the ground-truth class (i.e., detection or not), $\alpha_{t}$ the weighting factor to take into account data imbalance defined as $\alpha\in[0,1]$ for class 1 and $1-\alpha$ for class -1, and $\gamma>1$ is the focusing factor. This loss is especially interesting in radar data, since high RCS targets can be easily detected, but low RCS targets or targets located at a far distance are more challenging to detect, and this can be taken into account by the $\gamma$ parameter. In terms of training-testing split, 90% of the data from five of the seven recorded scenarios have been used to train the network using Adam optimizer, leaving 10% for validation. The network was trained using the DelftBlue Supercomputer [52] from TU Delft. The remaining two recorded scenarios are used as a test set, i.e., with data completely new, unseen for the network.
@@ VI Results
Two main performance metrics are used to evaluate the results of the proposed neural network: the usual probability of detection (Pd) and probability of false alarm (Pfa) metrics, and the Chamfer distance (CD). In both cases, the lidar data is used as a reference, either in the occupancy grid format for the Pd and Pfa computation, or in the point cloud format for the Chamfer distance. While different definitions are given for the Chamfer distance in the literature, in this work the following is used:
@@
  • No Doppler & Quantile: an approach similar to the one presented in [6], where only those power cells with values higher than the 0.9 quantile are kept and the rest is set to zero. Furthermore, the Doppler information is collapsed by taking the mean over the Doppler dimension. This is used to ’sparsify’ the data and speed up processing, with the risk of cutting out weakly reflecting targets. Since there is no Doppler data anymore, the DopplerEncoder subnetwork is removed from the general architecture of the proposed data-driven detector. It is important to note that the segmentation backbone has 13.2 million parameters, while the DopplerEncoder subnetwork only 76.9K. Therefore, the comparison between the full network and the network without the DopplerEncoder is possible without adding extra layers.
@@
  • OSCFAR: a 2D OS CFAR in range-angle, followed by a 1D OS CFAR in Doppler. While multiple different CFAR alternatives have been tested (i.e., different combinations of CA and OS CFAR detectors), only the best implementation is reported here for conciseness. An analysis with different variations has been presented in [18] for completeness. Following [53], the rank has been set to 0.75 times the number of training cells, and no guard cells have been used.
As it can be seen in Table III, the highest Pd and the lowest Chamfer distance is achieved by the proposed method while maintaining a similar Pfa. On the other hand, applying the quantile cut and removing the Doppler information similarly to [6] reduces the Pd from 62.13% to 52.97%, and worsens the Chamfer distance from 1.54m to 2.16m. Looking at the results for the other versions, it can be seen that this drop in performance is mostly due to the removal of the Doppler information. Using only the quantile-based threshold may be a good trade-off since the performance degradation is not substantial, but the computational cost is reduced. Looking at the version without the Temporal Coherence subnetwork, which is trained on single frames, it can be seen how all the metrics are worse than in the baseline. Thus, including temporal information in the network is a good strategy to boost performance, with the only downside of increasing slightly the training time due to the extra layers. Finally, it can be seen that the conventional OS CFAR is the method that performs the worst, with a much higher Chamfer distance of 6.73m.
In order to have a fairer comparison against the conventional CFAR detector, a 2D version of the proposed method has also been evaluated by disregarding the elevation information, as this can only be estimated rather poorly due to the unfavorable design of the radar array. To this end, the proposed NN has been trained without elevation information, discarding the virtual channels in the Z-direction and, thus, treating it as a ULA in the azimuth direction. For completeness, the implementation with quantile based threshold has also been assessed in this new analysis. The results are shown in Table III under the ’No Elevation cases’ rows. For these tests, the Pd of the OS CFAR approach is increased to 11.5%, but the Pfa is also raised. This is mainly due to detections triggered in the adjacent angle bins of a target generating ”ring like” patterns due to side lobes, a phenomenon also mentioned in [6]. Both the proposed method and the proposed method with the quantile based threshold are shown to outperform the conventional OSCFAR in the three metrics.
@@
Proposed Method 62.13 2.77 1.54
No Doppler & Quantile 52.97 2.63 2.16
No Doppler 50.44 2.50 2.13
Quantile 57.9 2.85 1.92
No Time (single frame) 58.08 2.63 2.16